Solve for y. 10x-9y=3
Answer by nerdybill(7384)

Solve for y. 10x-9y=3
.
10x-9y=3
-9y=-10x+3
9y=10x-3
y= 10x/9 - (3/9)
y= (10/9)x - 1/3